A rectangle is twice as long as it is wide. If its area is 50 sq. yd., find the length.

Found 2 solutions by JulietG, ewatrrr:
Answer by JulietG(1812)

L * W = Area
L = 2W
Substitute the value of L from the second equation into the first.
2W * W = 50
2W^2 = 50
Divide each side by 2
w^2 = 25
Take the square root of each side
w = 5
.
If the width is 5, then the length is twice that, or 10. 10*5=50
Success!

Answer by ewatrrr(24785)

 
Hi,
.
A = Lw
(2w)w = 50 yd^2
2w^2 = 50
w^2 = 25 (Tossing out the negative solution for unit measure)
w = 5yd and L = 10yd